# CyberLLMInstruct: A New Dataset for Analysing Safety of Fine-Tuned LLMs Using Cyber Security Data

### v2

@@ Abstract.
The integration of large language models (LLMs) into cyber security applications presents significant opportunities, such as enhancing threat analysis and malware detection, but can also introduce critical risks and safety concerns, including personal data leakage and automated generation of new malware. To address these challenges, we developed CyberLLMInstruct, a dataset of 54,928 instruction-response pairs spanning cyber security tasks such as malware analysis, phishing simulations, and zero-day vulnerabilities. The dataset was constructed through a multi-stage process. This involved sourcing data from multiple resources, filtering and structuring it into instruction-response pairs, and aligning it with real-world scenarios to enhance its applicability. Seven open-source LLMs were chosen to test the usefulness of CyberLLMInstruct: Phi 3 Mini 3.8B, Mistral 7B, Qwen 2.5 7B, Llama 3 8B, Llama 3.1 8B, Gemma 2 9B, and Llama 2 70B. In our primary example, we rigorously assess the safety of fine-tuned models using the OWASP top 10 framework, finding that fine-tuning reduces safety resilience across all tested LLMs and every adversarial attack (e.g. the security score of Llama 3.1 8B against prompt injection drops from 0.95 to 0.15). In our second example, we show that these same fine-tuned models can also achieve up to 92.50% accuracy on the CyberMetric benchmark. These findings highlight a trade-off between performance and safety, showing the importance of adversarial testing and further research into fine-tuning methodologies that can mitigate safety risks while still improving performance across diverse datasets and domains. The dataset creation pipeline, along with comprehensive documentation, examples, and resources for reproducing our results, is publicly available at https://github.com/Adelsamir01/CyberLLMInstruct.
conference: International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italy
@@ 1. Introduction
The increasing use of large language models (LLMs) in cyber security applications necessitates a rigorous examination of their benefits and potential safety risks. LLMs have shown exceptional capabilities in many text generation tasks, including code synthesis (Ságodi et al., 2024) and question answering (Raiaan et al., 2024), signalling their transformative potential across various tasks. However, this promise is accompanied by substantial safety risks, requiring focused attention from researchers and practitioners alike (Charan et al., 2023; Derner et al., 2023; Yao et al., 2024).
A crucial factor in the success and security of LLMs is the quality of their training data. Well-structured datasets enable models to generalise effectively across diverse cyber security scenarios, yet obtaining such datasets remains a significant challenge. Existing cyber security datasets can be broadly categorised into those designed for traditional machine learning, and those tailored for LLMs. Traditional datasets often fall short in size and diversity, limiting their capacity to support the intricate needs of LLMs, while many LLM-specific datasets remain inaccessible, hindering collaborative research and reproducibility. Furthermore, although realism is critical for accurately reflecting the complexities of real-world cyber security applications, counter-factual and synthetic data are equally important for preparing LLMs to address unknown and emerging threats (Xu et al., 2024; Ferrag et al., 2024b). Therefore, addressing these dataset limitations is crucial for ensuring that LLMs can reliably detect, prevent, and adapt to evolving cyber threats without being constrained by narrow training data or inaccessible resources.
These limitations highlight the urgent need for more robust, diverse, and realistic datasets to fine-tune LLMs for cyber security applications. This urgency is highlighted by the increasing misuse of generative AI tools like FraudGPT (Falade, 2023) and WormGPT (Firdhous et al., 2023) in cyberattacks, enabling adversaries to execute more sophisticated and scalable attacks. A recent study by Falade (2023) revealed how malicious LLMs can be exploited to generate phishing lures, impersonation schemes and deepfakes, amplifying the arsenal of cybercriminals and exposing significant vulnerabilities. Another example is the work done by Alotaibi et al. (2024), in which techniques like prompt engineering have coerced LLMs into generating malicious content, such as phishing emails and malware code. Other research (Roy et al., 2024) has shown that vanilla ChatGPT, without requiring advanced adversarial exploits, can create phishing websites that emulate trusted organisations and evade detection by anti-phishing systems. These findings demonstrate the critical need for datasets that prepare LLMs to handle both known and emerging threats in an evolving threat landscape.
Such datasets must encompass a mix of real-world attack data, security logs, and vulnerability reports alongside synthetic and counter-factual data to address known and unknown threats. The inclusion of synthetic data is essential for preparing LLMs to handle emerging attack strategies and novel scenarios. Robustness in this context refers to a dataset’s ability to provide diverse, comprehensive, and high-quality data that enhances an LLM’s ability to generalise across a range of cyber security tasks and adversarial conditions. Similarly, the robustness of an LLM relates to its ability to maintain reliable performance under varying conditions, including adversarial attacks, data gaps, and unknown threats.
This paper introduces CyberLLMInstruct, a novel dataset designed to enhance LLMs’ cyber security capabilities via fine-tuning. In addition to evaluating model behaviours before and after fine-tuning and analysing the safety risks associated with leveraging potentially misused cyber security data, we provide practical insights and guidelines for improving fine-tuning methods to balance performance and safety. CyberLLMInstruct fills a critical gap in existing datasets, which often lack the size, diversity, and practical relevance needed for robust evaluation and development of LLMs in cyber security contexts.
@@
Contributions. We make the following contributions in this work:
  • We release a dataset, CyberLLMInstruct, consisting of 54,928 instruction-response pairs that address cyber security tasks. The dataset was constructed through a systematic process, sourcing from authoritative cyber security resources, publicly available threat reports, and simulated scenarios. It also includes advanced topics such as malicious script generation, zero-day vulnerabilities, and adversarial examples for testing model robustness.
@@
  • Our two usage examples of CyberLLMInstruct led to key insights about the trade-offs between safety risks and performance gains via fine-tuning LLMs. For instance, while fine-tuning may enhance an LLM’s cyber security task performance significantly, it may also reduce its safety resilience. These insights call for more research to safeguard LLMs for cyber security applications.
Overall, this work establishes a strong foundation for advancing the secure deployment of LLMs in cyber security, while providing researchers with a critical resource to further explore the interplay between model performance and safety.
The rest of this paper is organised as follows. Section 2 provides an overview of related work, highlighting existing cyber security datasets and their limitations. Section 3 describes the CyberLLMInstruct dataset, including its creation process, design choices, and unique features. Section 4 presents two examples of using CyberLLMInstruct. Section 5 discusses the trade-offs between performance improvements and security vulnerabilities, the implications of our findings for real-world applications, and the limitations of our work. Section 6 concludes the paper with a summary of key insights and directions for future research.
@@ 2. Related Work
CyBERTuned (Jang et al., 2024) Large corpus for pretraining No No (text corpus) $\sim$700MB No direct vulnerability eval Pretraining LLMs for security awareness
CySecBERT (Bayer et al., 2024) Security news, CVE reports No No (text corpus) $\sim$4.3M documents Limited Domain-adaptive BERT for security tasks
SecQA (Liu, 2023) Multiple-choice Q&A No No (Q&A pairs) 127 Qs (v1) 115 Qs (v2) Not evaluated Basic security knowledge benchmarking
CyberMetric (Tihanyi et al., 2024) Large cyber security Q&A benchmark No No (Q&A format) 10,000 questions Minimal Evaluating LLM knowledge in cybersecurity
SVEN (He and Vechev, 2023) Secure vs. insecure code pairs Insecure code snippets No (code diffs) 803 fix pairs Some (prefix-tuning for safe vs. unsafe code) Code generation control (secure/insecure outputs)
CyberLLMInstruct Instruction-response cyber security dataset Yes (malicious + benign) Yes (full instruction format) 54,928 records Yes, tested with OWASP framework Fine-tuning LLMs, adversarial testing, security training
@@ 3. Resource Description
The CyberLLMInstruct dataset, comprising 54,928 records, is designed to cover diverse cyber security topics, formatted in a two-column structure containing instructions and responses. It encapsulates a broad spectrum of cyber security knowledge, including open-source intelligence, threat intelligence, attack techniques, and malware analysis. These categories were chosen based on their foundational importance in prior datasets (e.g., CyBERTuned (Jang et al., 2024), CySecBERT (Bayer et al., 2024)) and their relevance to real-world applications. Prompts guide LLMs through advanced scenarios, e.g., designing phishing campaigns, countering cyber security breaches, conducting social engineering simulations, exploring IoT security, network reconnaissance, and applications of such techniques. Unlike datasets that focus solely on defensive or descriptive cyber security knowledge, CyberLLMInstruct intentionally includes directly malicious content alongside benign discussions obtained mostly from educational resources such as Capture the Flag (CTF) challenges. This includes offensive security techniques, adversarial prompts, and examples of harmful code segments. For example, records related to malware do not merely describe malware threats but also include real-world attack simulations, malware demonstrations, and evasion techniques. CyberLLMInstruct dataset is designed for researchers and practitioners with intermediate to advanced knowledge in cyber security, particularly those working on LLM fine-tuning, adversarial testing, and cyber threat analysis. See Table 1 for a concise comparison of CyberLLMInstruct with related datasets.
[…]
  Malware (19,224)   Social Engineering (13,732)   DoS/DDoS (5,493)
  MITM (5,493)   Zero-Day (4,394)   Password (3,296)
  IoT (1,648)   Injection (1,648)
@@
Social engineering and phishing, representing 25% of the dataset, leverage human manipulation to compromise sensitive information. Denial-of-service (DoS, including distributed DoS – DDoS) attacks, and man-in-the-middle (MITM) attacks each account for 10% of the records, focusing on system disruption and eavesdropping, respectively. Zero-day exploits (8%) highlight the risks associated with unpatched vulnerabilities, while password attacks (6%) target weak or stolen credentials. Emerging threats such as Internet of Things (IoT) attacks (3%) and injection attacks (3%) showcase the vulnerabilities of interconnected devices and database systems.
Categories were assigned to the records using a semi-automatic methodology, combining manual (by the first author of the paper) and automated approaches utilising ChatGPT-4.0 and Gemini 1.5 Flash to ensure precision while optimising the efficiency of the process. The distribution of categories is detailed in Figure 1, showing how records were allocated in the dataset to align with the significance of each category in modern cyber security. For more details on how the categories were selected, please refer to Section 3.2.
@@ 3.1. Dataset Creation Process
The process began with Stage 1 “Data Collection”, where raw content was gathered from different sources covering real-world data sources. Authoritative resources, including NIST standards, research papers published in reputable venues, and industry reports (e.g., from the SANS Institute), were prioritised. Real-world incident reports, including data from the Common Vulnerabilities and Exposures (CVE) database, were also incorporated, along with hands-on examples from CTF challenges. Additionally, malicious cyber data, such as phishing campaigns, malware code, and known exploits, were included in the dataset to enhance the LLM’s effectiveness when confronted with malicious prompts. The full list of these resources and their justifications is provided with the dataset.
@@
In Stage 2 “Preliminary Filtering”, irrelevant materials for cyber security were removed after manual inspection by the first author.
@@
During Stage 3 “Data Structuring”, the dataset was formatted into a consistent two-column instruction-response format, transforming raw content into prompts and corresponding responses.
In Stage 4 “Domain Assignment”, data were categorised semi-automatically, distributing records across eight categories to comprehensively cover a broad range of cyber security topics. These categories included “Malware”, “Social Engineering and Phishing”, “DoS/DDoS Attacks”, “MITM Attacks”, “Zero-Day Exploits”, “Password Attacks”, “IoT Attacks”, and “Injection Attacks”, as outlined in Figure 1. This categorisation ensures that the dataset encompasses diverse and critical areas within the cyber security domain.
To maintain high quality, manual inspection was primarily conducted during Stage 5 “Manual Review”, where the paper’s first author verified the technical accuracy, relevance, and educational value of entries. Both benign and malicious entries were scrutinised to eliminate trivial or misleading content, ensuring alignment with current threats and best practices. A random sampling approach was used, with efforts made to balance the manual review across categories. Categories with higher percentages of entries received proportionally more attention to ensure comprehensive coverage. This selective yet systematic approach aimed to maximise reliability while managing the scope of the review, though some unreviewed records may carry a marginal risk of reduced accuracy.
In Stage 6 “Security Alignment”, adversarial considerations and red-teaming input were integrated to ensure that malicious examples were designed to realistically test the limits of the LLM’s compliance mechanisms. The focus was on exposing the LLM to adversarial and malicious prompts, alongside benign cyber security data, to challenge its ability to identify and respond to high-risk queries responsibly. This dataset was intentionally designed to push the boundaries of the LLM’s safety guardrails, recognising that in such scenarios, the LLM might fail or exhibit vulnerabilities, providing critical insights into its limitations.
Next, Stage 7 “AI-Assisted Tuning” used advanced LLMs (Gemini 1.5 Pro and GPT-4) to enhance language clarity, detail, and consistency. This refinement focused on improving the quality of both prompts and answers, making them more comprehensive and detailed. The process also addressed errors, reduced biases, and improved coherence across both malicious and benign examples. Following this stage, a final round of manual review was conducted for a random selected subset to ensure the outputs met the desired standards of precision, accuracy, and completeness.
Finally, Stage 8 “Final Integration and Packaging” consolidated all reviewed, categorised, and security-aligned records into a cohesive dataset. The result is a comprehensive resource designed to facilitate rigorous testing of LLMs, to enhance LLMs’ performance on cyber security tasks, and to challenge them with adversarial content.

### v3

# CyberLLMInstruct: A Pseudo-malicious Dataset Revealing Safety-performance Trade-offs in Cyber Security LLM Fine-tuning

@@ Abstract.
The integration of large language models (LLMs) into cyber security applications presents both opportunities and critical safety risks. We introduce CyberLLMInstruct, a dataset of 54,928 pseudo-malicious instruction-response pairs spanning cyber security tasks including malware analysis, phishing simulations, and zero-day vulnerabilities. Our comprehensive evaluation using seven open-source LLMs reveals a critical trade-off: while fine-tuning improves cyber security task performance (achieving up to 92.50% accuracy on CyberMetric), it severely compromises safety resilience across all tested models and attack vectors (e.g., Llama 3.1 8B’s security score against prompt injection drops from 0.95 to 0.15). The dataset incorporates diverse sources including CTF challenges, academic papers, industry reports, and CVE databases to ensure comprehensive coverage of cyber security domains. Our findings highlight the unique challenges of securing LLMs in adversarial domains and establish the critical need for developing fine-tuning methodologies that balance performance gains with safety preservation in security-sensitive domains.
large language models, cyber security, dataset, fine-tuning, adversarial testing, model safety, pseudo-malicious data
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 2025 Workshop on Artificial Intelligence and Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 Workshop on Artificial Intelligence and Security (AISec ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3733799.3762968isbn: 979-8-4007-1895-3/2025/10ccs: Computing methodologies Artificial intelligenceccs: Security and privacy Systems security
@@ 1. Introduction
The integration of large language models (LLMs) into cyber security applications presents both significant opportunities and critical safety risks (Charan et al., 2023; Derner et al., 2023; Yao et al., 2024). While LLMs show exceptional capabilities in tasks like code synthesis (Ságodi et al., 2024) and question answering (Raiaan et al., 2024), their application to cyber security domains requires careful examination of potential vulnerabilities.
A critical factor in LLM success is training data quality. Existing cyber security datasets often lack the size, diversity, and practical relevance needed for robust LLM fine-tuning (Xu et al., 2024; Ferrag et al., 2024b). Many datasets are either too narrow in scope or remain inaccessible, hindering reproducible research and effective model development.
The need for comprehensive cyber security datasets is further emphasised by the increasing misuse of generative AI tools like FraudGPT (Falade, 2023) and WormGPT (Firdhous et al., 2023), which enable sophisticated attacks including phishing campaigns, malware generation, and social engineering (Alotaibi et al., 2024; Roy et al., 2024). These developments highlight the critical need for datasets that prepare LLMs to handle both defensive applications and potential misuse scenarios.
This paper introduces CyberLLMInstruct [^1] https://github.com/Adelsamir01/CyberLLMInstruct – This repository contains all code and materials to reproduce the dataset used in the paper., a novel dataset designed to enhance LLMs’ cyber security capabilities via fine-tuning. CyberLLMInstruct consists of pseudo-malicious data, which contains instructions and descriptions of malicious cyber security actions without actual harmful code. Instead, it includes step-by-step descriptions and pseudo-code of how to perform these actions, such as malware creation, social engineering techniques, and various attack methodologies. This approach allows for comprehensive security testing while maintaining ethical boundaries. In addition to evaluating model behaviours before and after fine-tuning and analysing the safety risks associated with leveraging such cyber security data, we provide practical insights and guidelines for improving fine-tuning methods to balance performance and safety.
@@
We make the following contributions in this work:
  • We release a new dataset, CyberLLMInstruct, consisting of 54,928 pseudo-malicious instruction and response pairs that address cyber security tasks. The dataset was constructed through a systematic process, sourcing from authoritative cyber security resources, publicly available threat reports, and simulated scenarios. It includes step-by-step descriptions and pseudo-code for advanced cyber security topics such as malicious script generation, zero-day vulnerabilities, and adversarial examples for testing model robustness while maintaining ethical boundaries.
@@
  • Our two usage examples of CyberLLMInstruct led to key insights about the trade-offs between safety risks and performance gains via fine-tuning LLMs. For instance, while fine-tuning can enhance an LLM’s cyber security task performance significantly, it can also reduce its safety resilience. These insights call for more research to safeguard LLMs for cyber security applications.
Overall, this work establishes a strong foundation for advancing the secure deployment of LLMs for cyber security applications, while providing researchers with a critical resource to further explore the interplay between model performance and safety.
The rest of this paper is organised as follows. Section 2 provides an overview of related work, highlighting existing cyber security datasets and their limitations. Section 3 describes the CyberLLMInstruct dataset, including its creation process, design choices, and unique features. Section 4 presents two examples of using CyberLLMInstruct. Section 5 discusses the trade-offs between performance improvements and security vulnerabilities, the implications of our findings for real-world applications, and the limitations of our work with some future research direction. The last section concludes the paper.
@@ 2. Related Work
CyBERTuned (Jang et al., 2024)
Large corpus
for pretraining
No No (text corpus) $\sim$700MB
No direct
vulnerability eval
Pretraining LLMs for
security awareness
CySecBERT (Bayer et al., 2024)
Security news,
CVE reports
No No (text corpus)
$\sim$4.3M
documents
Limited
Domain-adaptive BERT
for security tasks
SecQA (Liu, 2023)
Multiple-choice
Q&A
No No (Q&A pairs)
127 Qs (v1)
115 Qs (v2)
Not evaluated
Basic security knowledge
benchmarking
CyberMetric (Tihanyi et al., 2024)
Large cyber security
Q&A benchmark
No No (Q&A format)
10,000
questions
Minimal
Evaluating LLM knowledge
in cybersecurity
SVEN (He and Vechev, 2023)
Secure vs. insecure
code pairs
Insecure code
snippets
No (code diffs)
803
fix pairs
Some (prefix-tuning for
safe vs. unsafe code)
Code generation control
(secure/insecure outputs)
CyberLLMInstruct
Instruction-response
cyber security
dataset
Yes
(malicious + benign)
Yes
(full instruction
format)
54,928
records
Yes, tested with
OWASP framework
Fine-tuning LLMs,
adversarial testing,
security training
@@ 3. Resource Description
The CyberLLMInstruct dataset, comprising 54,928 pseudo-malicious records, is designed to cover diverse cyber security topics, formatted in a two-column structure containing instructions and responses. The term pseudo-malicious refers to data that contains instructions and descriptions of malicious cyber security actions, but without actual harmful code. Instead, it includes step-by-step descriptions and pseudo-code of how to perform these actions, such as malware creation, social engineering techniques, and various attack methodologies. This approach allows for comprehensive security testing while maintaining ethical safeguards, as only pseudo-code and descriptive steps are included rather than functional malicious software, reducing the chance of direct misuse.
The dataset encapsulates a broad spectrum of cyber security knowledge, including open-source intelligence, threat intelligence, attack techniques, and malware analysis. These categories were chosen based on their foundational importance in prior datasets (e.g., CyBERTuned (Jang et al., 2024), CySecBERT (Bayer et al., 2024)) and their relevance to real-world applications. The dataset’s composition reflects real-world cyber threats, with malware-related content (35%), social engineering and phishing (25%), Denial-of-service (DoS, including distributed DoS – DDoS) attacks (10%), MITM attacks (10%), zero-day exploits (8%), password attacks (6%), and emerging threats like IoT and injection attacks (3% each). These categories not only capture prevalence but also illustrate different threat mechanisms: phishing exploits human manipulation, DoS and MITM target availability and confidentiality, zero-day exploits highlight risks of unpatched vulnerabilities, password attacks exploit weak credentials, while IoT and injection attacks expose weaknesses in connected devices and databases.
Prompts guide LLMs through advanced scenarios using pseudo-malicious content, including designing phishing campaigns, conducting social engineering simulations, exploring IoT security vulnerabilities, and network reconnaissance methodologies. Unlike datasets that focus solely on defensive or descriptive cyber security knowledge, CyberLLMInstruct intentionally includes pseudo-malicious instructional content alongside benign discussions obtained from educational resources such as Capture the Flag (CTF) challenges. For example, records related to malware include step-by-step descriptions of attack simulations, conceptual malware demonstrations, and evasion technique explanations without providing actual executable malicious code.
CyberLLMInstruct dataset is designed for researchers and practitioners with intermediate to advanced knowledge in cyber security, particularly those working on LLM fine-tuning, adversarial testing, and cyber threat analysis. See Table 1 for a concise comparison of CyberLLMInstruct with related datasets.
@@
Categories were assigned to the records using a semi-automatic methodology, combining manual (by the first author of the paper) and automated approaches utilising GPT-4 and Gemini 1.5 Flash to ensure precision while optimising the efficiency of the process. The distribution of categories is detailed in Figure 1, showing how records were allocated in the dataset to align with the significance of each category in modern cyber security. For more details on how the categories were selected, please refer to Section 3.4.
@@ 3.1. Dataset Creation Process
The process began with Stage 1 “Data Collection”, where raw content was gathered from different sources covering real-world data sources. Authoritative resources, including NIST standards, research papers published in reputable venues, and industry reports (e.g., from the SANS Institute), were prioritised. Real-world incident reports, including data from the Common Vulnerabilities and Exposures (CVE) database, were also incorporated, along with hands-on examples from CTF challenges. Additionally, pseudo-malicious cyber security data, such as educational descriptions of phishing campaign structures, malware analysis methodologies, and exploit technique explanations, were included in the dataset to enhance the LLM’s effectiveness when confronted with security-related prompts while maintaining ethical boundaries through instructional rather than executable content. A comprehensive list of all data sources used in the collection process is provided in Appendix A.
3.2. Data Processing Stages
Following data collection, CyberLLMInstruct underwent systematic processing through seven additional stages:
@@ 3.2. Data Processing Stages
Stage 2 – Preliminary Filtering: Irrelevant materials for cyber security were removed after manual inspection by the first author, eliminating content that did not contribute to security knowledge or skills.
@@
Stage 3 – Data Structuring: Raw content was formatted into a consistent two-column instruction-response structure, transforming diverse source materials into standardised training prompts and corresponding detailed responses.
Stage 4 – Domain Assignment: Data were categorised semi-automatically across eight security domains using a combination of manual review and automated classification with GPT-4 and Gemini 1.5 Flash, ensuring comprehensive coverage of critical cyber security areas.
Stage 5 – Manual Review: Technical accuracy, relevance, and educational value were verified through systematic manual inspection. Both benign and pseudo-malicious entries were scrutinised to eliminate trivial or misleading content while ensuring ethical boundaries were maintained, with proportional attention given to larger categories.
Stage 6 – Security Alignment: Adversarial considerations were integrated to ensure pseudo-malicious examples realistically test LLM compliance mechanisms while maintaining ethical boundaries. This stage focused on exposing models to challenging educational scenarios that push the boundaries of safety guardrails without providing actual harmful executable content.
Stage 7 – AI-Assisted Tuning: Advanced LLMs (Gemini 1.5 Pro and GPT-4) enhanced language clarity, detail, and consistency across all entries, improving both prompt quality and response comprehensiveness while maintaining technical accuracy and ethical boundaries for pseudo-malicious content.
Stage 8 – Final Integration: All processed, reviewed, and aligned records were consolidated into the final dataset, creating a comprehensive resource for LLM fine-tuning.
3.3. Data Source Breakdown
CyberLLMInstruct incorporates content from diverse sources to ensure comprehensive coverage:
  • Capture The Flag (CTF) Challenges (27%): Practical security puzzles covering web exploitation, cryptography, reverse engineering, and forensics from platforms like PicoCTF, OverTheWire, and VulnHub.
  • Academic Research Papers (22%): Peer-reviewed publications from security conferences (CCS, USENIX Security, NDSS) covering emerging threats, novel attack techniques, and defense mechanisms.
  • Industry Reports (18%): Threat intelligence reports from organisations like MITRE, NIST, and security vendors documenting real-world attack campaigns and IoCs.
  • CVE Database (15%): Detailed vulnerability descriptions, exploit techniques, and remediation guidance from the National Vulnerability Database.
  • Security Training Materials (10%): Content from certified ethical hacking courses, penetration testing methodologies, and security awareness training.
  • Malware Samples and Analysis (8%): Reverse engineering reports, malware family analysis, and behavioural descriptions from security research teams.